OpenStudy (donnie1999):

Celeste wants to have her hair cut and permed and also go to lunch. She knows she will need $86. The perm costs twice as much as her haircut and she needs $5 for lunch. How much does the perm cost?

OpenStudy (donnie1999):

@crissyyyxoxo

OpenStudy (donnie1999):

@aloud

OpenStudy (donnie1999):

@Y.R.N.

OpenStudy (anonymous):

$40.50

OpenStudy (anonymous):

and ill exsplane y

OpenStudy (anonymous):

86-5=81 81/2=4.5

OpenStudy (donnie1999):

it was wrong :(

OpenStudy (anonymous):

Let x = cost of haircut then 2x = cost of perm . x + 2x + 5 = 59 3x + 5 = 59 3x = 54 x = $18 (haircut) . Perm: 2x = 2(18) = $36 . Lunch: $5